OpenStudy (anonymous):

I WILLL MEDAL AND FAN!!! An expression is shown below: f(x) = -16x^2 + 22x + 3 Part A: What are the x-intercepts of the graph of the f(x)? Show your work. Part B: Is the vertex of the graph of f(x) going to be a maximum or minimum? What are the coordinates of the vertex? Justify your answers and show your work. Part C: What are the steps you would use to graph f(x)? Justify that you can use the answers obtained in Part A and Part B to draw the graph.

OpenStudy (anonymous):

part A, set \[-16x^2 + 22x + 3 =0\] and solve for \(x\)

OpenStudy (anonymous):

it is not that straight forward, probably have to use the quadratic formula i would start with \[16x^2-22x-3=0\] and that one

OpenStudy (anonymous):

Quadratic formula: \[x = \frac{ -b \sqrt{b^{2}-4ac}}{2a }\]

OpenStudy (anonymous):

\[x = \frac{ -b\color{red}{\pm} \sqrt{b^{2}-4ac}}{2a }\]

OpenStudy (campbell_st):

just multiply -16 by 3 = -48 find the factors that add to 22...... 24 - 2 then its (-16x + 24)(-16x - 2) ------------------- -16 factoring and you get 8(-2x + 3)-2(8x + 1) ------------------ -16 cancel the common factors (-2x +3)(8x + 1) = 0 so solve -2x + 3 = 0 and 8x + 1 = 0 for the intercepts

OpenStudy (campbell_st):

so for the man or min... look at the sign of the leading term....

OpenStudy (anonymous):

So Is It A Maximum?

OpenStudy (campbell_st):

thats correct...

OpenStudy (campbell_st):

the vertex.... so it can be found by 1st finding the line of symmetry in the general form y = ax^2 + bx + c the line of symmetry is \[x = \frac{-b}{2a}\] in your question b = 22 and a = -16 the vertex is on the line of symmetry... so substitute the value of x into the original equation to find the vertex... then you can use A and B to graph the equation... you know x - intercepts, vertex, concavity and by letting x = 0 you can find the y- intercept... hope it helps

OpenStudy (anonymous):

Thank You So Much Campbell!